Find the magnitude and direction of the vector represented by the following pairs of components:

(a) Ax = -8.60 cm, Ay = 5.20 cm; (b) Ax = -9.70 m., Ay = -2.45 m; (c) Ax = 7.75 km, Ay = -2.70 km.


1
Expert's answer
2021-12-16T11:33:33-0500

(a) "A=\sqrt{A_x^2+A_y^2}=10\text{ cm},\theta=\arctan(A_x/A_y)=149°."

(b) "A=\sqrt{A_x^2+A_y^2}=10\text{ cm},\theta=\arctan(A_x/A_y)=194°."

(c) "A=\sqrt{A_x^2+A_y^2}=8.2\text{ cm},\theta=\arctan(A_y/A_x)=-19°."

All angles are from positive x-axis.